An apparatus for determining molar mass of unknown gases includes a large sealed glass vessel under vacuum on a balance held at 0.0 ⁰C. To calibrate the volume of the vessel, nitrogen gas was added until the pressure reached 120.0 kPa. The container mass increased by 25.0 g. What is the volume of the vessel (L)? N2 gas has a molar mass of 14.01 g/mol.
1
Expert's answer
2020-10-06T14:11:32-0400

As we know from Ideal gas equation that "PV =nRT" .............(1)

Now in our question n = number of mole = "25\/14.01 = 1.7844"

, P= "120 * 10^3 Pa" ...... T "= 0+273 = 273K" ,R="8.314*10^3 L.Pa.K^(-1).mol^(-1)" Now from equation 1.... Volume "= V =1.7844*8.3148*(10^3)*273\/{120*10^3}"=33.75 Litre hence "Volume= 33.75 L" Answer.